# Entanglement Hamiltonians and the quasiparticle picture

### v1

@@ Abstract
The entanglement Hamiltonian (EH) provides the most comprehensive characterization of bipartite entanglement in many-body quantum systems. Ground states of local Hamiltonians inherit this locality, resulting in EHs that are dominated by local, few-body terms. Unfortunately, in non-equilibrium situations, analytic results are rare and largely confined to continuous field theories, which fail to accurately describe microscopic models. To address this gap, we present an exact analytic result for the EH following a generic quantum quench in non-interacting fermionic models. This derivation adapts the celebrated quasiparticle picture to the EH, providing detailed insights into its physical properties. The resulting analytic formula serves as a foundation for engineering EHs in quantum optics experiments.
@@ Introduction.
It is thus very natural to explore and search for other physical situations where the EH is local. To date, local EHs have primarily been derived in the realm of conformal field theory (CFT), where conformal transformations enable the mapping of ground state results to more complex scenarios [29, 30, 31, 32], such as quantum quenches [32, 33]. Notwithstanding some studies on EHs after quantum quenches in lattice models [34, 35], it is still unclear to what extent the locality of the EH is preserved in the absence of conformal symmetry. The goal of this Letter is to initiate a systematic study of the EHs in non-equilibrium situations. We focus here on free fermionic models and show that in general the EH after a global quantum quench can be reconstructed from the quasiparticle picture (QPP) [36, 37, 38, 39]. Originally developed to describe entanglement entropy growth after a quantum quench in integrable systems, this picture posits that entanglement is carried by pairs of quasiparticles emitted from the initial state. As these quasiparticles propagate through the system, they spread entanglement in a manner that can be quantitatively tracked and predicted. It has since been shown to be applicable to the calculation of other quantities such as some correlation functions [40], negativity [41, 42], full counting statistics [43, 44, 45], symmetry resolved entanglement [45, 46, 47], operator entanglement [48, 49], and the entanglement asymmetry [50, 51, 52]. However, despite the fact that it contains information about most of these quantities, the application of the QPP to describe the EH has remained extremely elusive. In this letter, we fill this void and derive a very compact form for the EH after a generic integrable quench in a free-fermionic model.
@@
Here the pure RDM comes from the pairs of quasiparticles that are both in $A$, while the entangling one, with EH $K_{A,\text{ QP}}^{(t)}$, describes the pairs shared between $A$ and $\bar{A}$, which are entirely responsible for the entanglement. The main result of this paper is that $K_{A,\text{ QP}}^{(t)}$ takes the two-body form
@@
where $n(k)=\operatorname{Tr}[\rho\,c^{\dagger}_{k}c_{k}]$. As a conserved quantity, $\eta(k)$ can be computed in the initial state without solving the dynamics. Therefore, this form of the EH clearly enables its reconstruction without needing to solve the dynamics. We stress that the main property of the form (3) is to be entirely determined by two-body terms with a kernel that depend on the distance $z$ in a light-cone fashion, see also Fig. 1 for an explicit example.
@@
The majority of the rest of this Letter is devoted to deriving this result and presenting several checks, numerical and analytical, of its validity. Before this however, we make some brief comments on the structure of (3). Immediately after the quench, the RDM is entirely captured by $\rho_{\text{pure}}^{(t)}$ since no quasiparticle are yet shared. As time grows, the quasiparticles pass through the entangling points moving between $A$ and $\bar{A}$. Accordingly, the pure part $\rho_{\text{pure}}^{(t)}$ becomes progressively smaller, while the entangling one grows. The entangling part is governed by the kernels $\mathcal{K}_{L,R}(x,z)$ which can be interpreted as the contributions from the left $(L)$ and right $(R)$ moving quasiparticles which spread from the left and right edges of $A$ respectively. From the structure of Eqs. 4 and 5, we see that for $t<\ell/(4v_{\rm max})$, where $v_{\rm max}=\max(v_{k})$, the growth happens inside light-cones centered in 00 and $\ell$, as expected from the QPP. The real space structure of the EH is determined by the initial state through $\eta(k)$.
At sufficiently large times after the quench, the system relaxes locally to a stationary state which, generically for a free model is a generalized Gibbs ensemble (GGE) that incorporates all conserved charges. In the RDM, (2), the pure part no longer contributes, and the EH (3) fully characterizes the RDM. In this limit, the time dependence of $\mathcal{K}_{L,R}$ drops out and asymptotic value of the EH (3) is
$=\int\frac{\mathrm{d}k}{2\pi}\,\eta(k)\,c_{k}^{\dagger}c_{k},$
This light-cone behavior and eventual relaxation is especially evident when the quenching Hamiltonian has a linear dispersion, $\varepsilon_{k}=vk$. In this case the quasiparticle velocity $v$ is independent of $k$ and the Heaviside functions can be taken outside of the mode integrals in (4), (5). Moreover, if $\eta(k)=\beta\varepsilon_{k}$, as happens after a quench in CFT [54], we find that
where $T$ and $\overline{T}$ are the right and left moving stress-energy operators. Eq. (8) agrees with previous results for the post-quench entanglement Hamiltonian in a CFT [32, 33]. Note that in this case the system relaxes locally to a Gibbs state in which only the Hamiltonian appears [55] and accordingly the EH has only short ranged hopping terms. This however, is not generically the case and $K_{A,\text{ QP}}^{(t)}$ can have a more intricate structure depending on both the initial state as encoded in $\eta(k)$ and $v_{k}$.
@@
In Figure 1 we plot the values of the nearest neighbor and beyond nearest neighbor terms of $K_{A,\text{ QP}}^{(t)}$ for a particular case (see below for details). What is evident, is that each of the terms exhibits a light cone like spreading emitting from the subsystem edges. Moreover we see that the quasiparticle structure has endowed the hopping terms with a parity effect such that at finite times hopping over an odd number of sites is even with respect to reflection about the subsystem center while hopping over an even number of sites is odd with respect to this. As a result, in the long time limit only odd site hopping terms remain. From Figure 2 we also see that the relative strength of these hopping terms decays with a power law behavior as a function of hopping distance. Overall it is remarkable that the complicated shapes of the EH reported in the figures are captured perfectly by a simple quasiparticle form as Eq. 3.
Figure 1: EH couplings $h_{j,j+z}$ between fermions at distances $z=1,2,3,4$ after a quench from the dimer state. We study an interval of length $\ell=800$ at different times $t$ after the quench. The symbols are obtained from Eq. 24 using the correlation matrix (22), while the dashed black line is the QPP prediction (3). The couplings are real for odd distances $z$, while they are imaginary for even $z$. For $t\to\infty$, (gray crosses), the data match the GGE, cf. Eq. 7 (red dashed line). We observe a perfect agreement, with minor deviations near the endpoints where the QPP is not expected to work. For times $t/\ell\leqslant 0.25$ (blue and orange symbols) the couplings are non zero only inside of two light-cones centered at the endpoints. For $t>0.25\ell$, the two light-cones merge and the EH relaxes toward the GGE.
@@
As a starting point of our computation, we assume that at $t=0$ the initial state is a squeezed state of the form
|$\begin{split}\ket{\psi^{(0)}}&=\exp\!\left\{\mathrm{i}\int_{k>0}\frac{\mathrm{d}k}{2\pi}\,\mathcal{M}(k)\,c^{\dagger}_{k}c^{\dagger}_{-k}\right\}\ket{0}\,,\end{split}$ (9)
where |$\ket{0}$ is the vacuum state of the post-quench Hamiltonian and $\mathcal{M}(k)$ is some odd function whose particular form is unimportant. In states of the form (9), the population $n(k)$ of the fermionic modes is
@@
An important property of the state (9) is that it has a natural pair structure, in which only modes of opposite momenta are entangled. For this reason, it would seem that Eq. 9 is a very specific choice, however squeezed states appear ubiquitously in quantum quenches in integrable models, see, e.g., Ref. [56].
To show that the RDM follows a QPP, in the spirit of Ref. [57] we employ an hydrodynamic description. We introduce hydrodynamic cells of length $\Delta$ much larger than the lattice spacing $a$ (or any other short distance cutoff) but smaller than the typical length $\ell$ of our system $a\ll\Delta\ll\ell$, and we impose periodic boundary conditions at the edges of the cell. Quasiparticles are defined as wave-packets localized inside an hydrodynamic cell, obtained by performing a Fourier transform only inside the cell [57]
@@
$\rho_{x_{0},k}^{(0)}=n(k)b^{\dagger}_{x_{0},k}b_{x_{0},k}b^{\dagger}_{x_{0},-k}b_{x_{0},-k}\\ +(1-n(k))(1-b^{\dagger}_{x_{0},k}b_{x_{0},k}b^{\dagger}_{x_{0},-k}b_{x_{0},-k})\\ +\mathrm{i}\sqrt{n(k)(1-n(k))}(b^{\dagger}_{x_{0},k}b^{\dagger}_{x_{0},-k}-b_{x_{0},-k}b_{x_{0},k})\,.$ (13)
In this approximation, only pairs of quasiparticles with opposite momenta and occupying the same cell are entangled with each others [57].
@@
$\rho^{(t)}_{x_{0},k}=n(k)\,b^{\dagger}_{x_{0}+v_{k}t,k}b_{x_{0}+v_{k}t,k}b^{\dagger}_{x_{0}-v_{k}t,-k}b_{x_{0}-v_{k}t,-k}\\ +(1-n(k))(1-b^{\dagger}_{x_{0}+v_{k}t,k}b_{x_{0}+v_{k}t,k}b^{\dagger}_{x_{0}-v_{k}t,-k}b_{x_{0}-v_{k}t,-k})\\ +\mathrm{i}\sqrt{n(k)(1-n(k))}(b^{\dagger}_{x_{0}+v_{k}t,k}b^{\dagger}_{x_{0}-v_{k}t,-k}\\ -b_{x_{0}-v_{k}t,-k}b_{x_{0}+v_{k}t,k})\,.$ (15)
@@
$\rho_{A,x_{0},k}^{(t)}=n(k)\,b^{\dagger}_{x_{0}+v_{k}t,k}b_{x_{0}+v_{k}t,k}\\ +(1-n(k))(1-b^{\dagger}_{x_{0}+v_{k}t,k}b_{x_{0}+v_{k}t,k})\,.$ (16)
By following the trajectories of the quasiparticles, we see that at time $t$ the quasiparticles that are shared between $A$ and its complement are the right-movers that originated at $x_{0}\in[-v_{k}t,\min(v_{k}t,\ell-v_{k}t)]$ and the left-movers that started at $x_{0}\in[\ell-\min(\left|v_{k}\right|t,\ell-\left|v_{k}\right|t),\ell+\left|v_{k}\right|t]$. Expressing everything in terms of the current position of the quasiparticle $x=x_{0}+v_{k}t$ for right-movers, $x=x_{0}-\left|v_{k}\right|t$ for left movers, we have
@@
To find the entanglement Hamiltonian of the mixed part of Eq. 2, we rewrite Eq. 16 as an exponential, using the property $\left(b^{\dagger}b\right)^{2}=b^{\dagger}b$. We obtain $\rho_{A}^{(t)}=e^{-K_{A,\text{ QP}}^{(t)}}/Z_{A}$, where $Z_{A}$ is a normalization and
@@
$S^{(\alpha)}_{A}(t)=\int\frac{dk}{2\pi}\,\min(2|v_{k}|t,\ell)\,h_{\alpha}(n(k))\,,$ (21)
@@
where $J_{\nu}(z)$ is Bessel’s function and we have introduced the asymptotic value of the correlation matrix $C_{i,j}^{(\infty)}=\delta_{i,j}/4+(\delta_{i,j-1}+\delta_{i,j+1})/2$.
Figure 2: EH couplings $h_{j,j+z}$ after a quench from the dimer state. We consider an interval of length $\ell=800$ at time $t/\ell=0.2$. In the main plot we report the couplings (obtained using Eqs. 24 and 22) for different (odd) distances $z$. The black dashed line is the prediction (3). The couplings present an oscillatory behavior with a peak which is slowly decreasing as the distance $z$ increases. The inset reports the peak value as a function of the separation $z$ in logarithmic scales, for $z$ odd from $3$ to $17$. The decay of the peak is compatible with a power law with power $\alpha=-0.6$ (black dotted line).
From the correlation matrix $C_{A}$ restricted to the subsystem $A$ it is possible to directly compute the entanglement Hamiltonian [59, 60, 61, 62]. Writing the lattice entanglement Hamiltonian as
@@
In order to compare the result of Eq. 24 with our prediction in Eq. 3, however, we need to take into account that at finite time Eq. 3 only describes the low lying part of the entanglement spectrum. Since according to Eq. 24 the higher part of the entanglement spectrum is given by the eigenvalues of $C_{A}$ that are close to either 00 or $1$, we introduce a cut-off on the spectrum, projecting out all the eigenvalues of Eq. 22 which are smaller than $10^{-4}$ or larger than $1-10^{-4}$.
@@
Conclusions and Outlook. In this Letter, we derived a quasiparticle prediction, Eq. 3, for the EH after a global quantum quench in any system of non-interacting fermions. Remarkably, this simple form effectively captures the complex structure of the EH.
Our result paves the way for many intriguing generalizations. One key advantage of the QPP is its adaptability for calculating the entanglement entropy of disjoint intervals [36, 63], as well as more complex quantities like the negativity [41, 42]. Generalizing these results to an operator level, such as calculating the entanglement Hamiltonian of disjoint intervals or the negativity Hamiltonian [64, 65, 66], would be extremely interesting. Additionally, the QPP applies also for certain dissipative systems [67, 68, 69], where also we could think of adapting our derivation. Furthermore, the QPP has been recently combined with dimensional reduction [70, 71] to describe the entanglement entropy [72] and asymmetry [73] in higher dimensional non-interacting models. The same could be done also for the EH. A significant challenge is generalizing our result to interacting integrable models, where the QPP breaks down in the calculation of Rényi entanglement entropy [74] and charged moments [45].
@@
Acknowledgements.
We thank F. Ares, L. Capizzi, O. Castro-Alvaredo, B. Doyon, G. Di Giulio, J. Dubail, S. Murciano, and P. Zoller for fruitful discussions. All authors acknowledge support from ERC under Consolidator Grant number 771536 (NEMO). FR acknowledges support from the project ‘Artificially devised many-body quantum dynamics in low dimensions - ManyQLowD’ funded by the MIUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 - Grant 2022R35ZBF.

### v2

@@ Abstract
The entanglement Hamiltonian (EH) provides the most comprehensive characterization of bipartite entanglement in many-body quantum systems. Ground states of local Hamiltonians inherit this locality, resulting in EHs that are dominated by local, few-body terms. Unfortunately, in non-equilibrium situations, analytic results are rare and largely confined to continuous field theories, which fail to accurately describe microscopic models. To address this gap, we present an analytic result for the EH following a quantum quench in non-interacting fermionic models, valid in the ballistic scaling regime. The derivation adapts the celebrated quasiparticle picture to operators, providing detailed insights into its physical properties. The resulting analytic formula serves as a foundation for engineering EHs in quantum optics experiments.
@@ Introduction.
It is thus very natural to explore and search for other physical situations where the EH is local. To date, local EHs have primarily been derived in the realm of conformal field theory (CFT), where conformal transformations enable the mapping of ground state results to more complex scenarios [29, 30, 31, 32], such as quantum quenches [32, 33]. Notwithstanding some studies on EHs after quantum quenches in lattice models [34, 35], it is still unclear to what extent the locality of the EH is preserved in the absence of conformal symmetry. The goal of this Letter is to initiate a systematic study of the EHs in non-equilibrium situations. We focus here on free fermionic models and show that in general the EH after a global quantum quench can be reconstructed from the quasiparticle picture (QPP) [36, 37, 38, 39]. Originally developed to describe entanglement entropy growth after a quantum quench in integrable systems, this picture posits that entanglement is carried by pairs of quasiparticles emitted from the initial state. As these quasiparticles propagate through the system, they spread entanglement in a manner that can be quantitatively tracked and predicted. It has since been shown to be applicable to the calculation of other quantities such as some correlation functions [40], negativity [41, 42], full counting statistics [43, 44, 45], symmetry resolved entanglement [45, 46, 47], operator entanglement [48, 49], and the entanglement asymmetry [50, 51, 52]. However, despite the fact that it contains information about most of these quantities, the application of the QPP to describe the EH has remained extremely elusive. In particular, while the Rényi entropies indicate that the spectrum of the RDM admits a QPP it is far from trivial that the matrix itself does. In this letter, we fill this void and derive a very compact form for the EH after a generic integrable quench in a free-fermionic model. We proceed by first presenting the result, followed by its derivation and then present some analytic and numerical checks of its veracity.
@@
Figure 1: EH couplings $h_{j,j+z}$ between fermions at distances $z=1,2,3,4$ after a quench from the dimer state. We study an interval of length $\ell=800$ at different times $t$ after the quench. The symbols are obtained from Eq. 24 using the correlation matrix (22), while the dashed black lines are the kernels $\mathcal{K}_{R}(j/\ell,z)+\mathcal{K}_{L}(j/\ell,z)$ of the QPP prediction (3) for different distances $z$. The couplings are real for odd distances $z$, while they are imaginary for even $z$. For $t\to\infty$, (gray crosses), the data match the GGE, cf. Eq. 19 (red dashed line). We observe a perfect agreement, with minor deviations near the endpoints where the QPP is not expected to work. For times $t/\ell\leqslant 0.25$ (blue and orange symbols) the couplings are non zero only inside of two light-cones centered at the endpoints. For $t>0.25\ell$, the two light-cones merge and the EH relaxes toward the GGE.
@@
Here, the tensor product is inherited from the quasiparticle structure of the Hilbert space in the following way. For every quasiparticle residing within $A$ we check if its correlated partner is also inside $A$ or not. If it is, then this pair does not generate entanglement and contributes to $\rho_{\rm pure}^{(t)}$. If it is not, then this generates entanglement and the quasiparticle contributes to $K_{A,\text{ QP}}^{(t)}$. The main result of this paper is that in the scaling limit of large times $t$ and large subsystem sizes $\ell$, with fixed ratio $\ell/t$, the EH $K_{A,\text{ QP}}^{(t)}$ takes the quadratic form
@@
where $n(k)=\operatorname{Tr}[\rho\,c^{\dagger}_{k}c_{k}]$. As a conserved quantity, $\eta(k)$ can be computed in the initial state without solving the dynamics. Therefore, this form of the EH clearly enables its reconstruction without needing to solve the dynamics. We stress that the main property of the form (3) is to be entirely determined by quadratic terms with a kernel that depends on the distance $z$ in a light-cone fashion, see also Fig. 1 for an explicit example.
@@
The majority of the rest of this Letter is devoted to deriving this result and presenting several checks, numerical and analytical, of its validity. Before this however, we make some brief comments on the structure of (3). Immediately after the quench, no quasiparticles are yet shared between $A$ and its complement. Accordingly, the full Hilbert space is identified with the one describing pairs, both of which are inside $A$, while the other one is empty, and the RDM is entirely captured by $\rho_{\text{pure}}^{(t)}$. As time grows, the quasiparticles pass through the entangling points moving between $A$ and $\bar{A}$. As a result, the dimension of the Hilbert space of the shared pairs increases while that of the unshared quasiparticles decreases. At long times, the Hilbert space of the shared quasiparticles becomes that of the full subsystem while that of the unshared quasiparticles becomes the empty set. The entangling part is governed by the kernels $\mathcal{K}_{L,R}(x,z)$ which can be interpreted as the contributions from the left $(L)$ and right $(R)$ moving quasiparticles which spread from the left and right edges of $A$ respectively. From the structure of Eqs. 4 and 5, we see that for $t<\ell/(4v_{\rm max})$, where $v_{\rm max}=\max(v_{k})$, the growth happens inside light-cones centered in 00 and $\ell$, as expected from the QPP. The real space structure of the EH is determined by the initial state through $\eta(k)$.
@@
In Fig. 1 we plot the values of the nearest neighbor and beyond nearest neighbor terms of the entanglement Hamiltonian, $\mathcal{K}_{L}(x,z)+\mathcal{K}_{R}(x,z),\,z=1,\dots,4$ for a particular case (see below for details). What is evident, is that each of the terms exhibits a light cone like spreading emitting from the subsystem edges. Moreover we see that the quasiparticle structure has endowed the hopping terms with a parity effect such that at finite times hopping over an odd number of sites is even with respect to reflection about the subsystem center and real, while hopping over an even number of sites is odd with respect to this and imaginary. In the long time limit only odd site hopping terms remain.
@@
As a starting point of our computation, we assume that at $t=0$ the initial state is a squeezed state of the form [58]
|$\begin{split}\ket{\psi^{(0)}}&=\mathcal{N}\exp\!\left\{\mathrm{i}\int_{k>0}\frac{\mathrm{d}k}{2\pi}\,\mathcal{M}(k)\,c^{\dagger}_{k}c^{\dagger}_{-k}\right\}\ket{0}\,,\end{split}$ (7)
where |$\ket{0}$ is the vacuum state of the post-quench Hamiltonian, $\mathcal{N}$ is the normalization and $\mathcal{M}(k)$ is some odd function whose particular form is unimportant. In states of the form (7), the population $n(k)$ of the fermionic modes is
@@
The choice of a state of the form (7) greatly simplifies the derivation, since its pair structure, in which only modes of opposite momenta are entangled, can be read directly from Eq. 7. Nevertheless, the final result is more general than the stated derivation, describing all free-fermionic quenches in which the initial state leads to the emission of entangled pairs of quasiparticles. Moreover it can be straightforwardly extended to account for states which have a more intricate quasiparticle structure [53]. As an example, in the next section we will study numerically a quench from the dimer state, whose scaling behavior, as can be seen from Figs. 1 and 2, is faithfully captured by our prediction, despite the initial state not being of the form (7).
Figure 2: EH couplings $h_{j,j+z}$ after a quench from the dimer state. We consider an interval of length $\ell=800$ at time $t/\ell=0.2$. In the main plot we report the couplings (obtained using Eqs. 24 and 22) for different (odd) distances $z$. The black dashed lines are the kernels $\mathcal{K}_{R}(j/\ell,z)+\mathcal{K}_{L}(j/\ell,z)$ of the prediction (3). The couplings present an oscillatory behavior with a peak which is slowly decreasing as the distance $z$ increases. The inset reports the couplings $h_{j,j+z}^{(\infty)}$ in the GGE which are approached at large times, for the same value of $z$ in the main plot. We see that the EH at intermediate time has a significantly different structure from the asymptotic value, demonstrating that the large time limit is reached in a highly non-trivial way.
To show that the RDM follows a QPP, in the spirit of Ref. [54] we employ a hydrodynamic description. We introduce hydrodynamic cells of length $\Delta$ much larger than the lattice spacing $a$ (or any other short distance cutoff) but smaller than the typical length $\ell$ of our system $a\ll\Delta\ll\ell$, and we impose periodic boundary conditions at the edges of the cell. Quasiparticles are defined as wave-packets localized inside a hydrodynamic cell, obtained by performing a Fourier transform only inside the cell [54]
@@
$\rho_{x_{0},k}^{(0)}=n(k)\hat{n}_{x_{0},k}\hat{n}_{x_{0},-k}\\ +(1-n(k))(1-\hat{n}_{x_{0},k})(1-\hat{n}_{x_{0},-k})\\ +\mathrm{i}\sqrt{n(k)(1-n(k))}(b^{\dagger}_{x_{0},k}b^{\dagger}_{x_{0},-k}-b_{x_{0},-k}b_{x_{0},k})\,.$ (11)
where $\hat{n}_{x_{0},k}=b^{\dagger}_{x_{0},k}b_{x_{0},k}$. In this approximation, only pairs of quasiparticles with opposite momenta and occupying the same cell are entangled with each others [54].
@@
$\rho^{(t)}_{x_{0},k}=n(k)\,\hat{n}_{x_{0}+v_{k}t,k}\hat{n}_{x_{0}-v_{k}t,-k}\\ +(1-n(k))(1-\hat{n}_{x_{0}+v_{k}t,k})(1-\hat{n}_{x_{0}-v_{k}t,-k})\\ +\mathrm{i}\sqrt{n(k)(1-n(k))}(b^{\dagger}_{x_{0}+v_{k}t,k}b^{\dagger}_{x_{0}-v_{k}t,-k}\\ -b_{x_{0}-v_{k}t,-k}b_{x_{0}+v_{k}t,k})\,.$ (13)
@@
$\rho_{A,x_{0},k}^{(t)}=n(k)\,\hat{n}_{x_{0}+v_{k}t,k}+(1-n(k))(1-\hat{n}_{x_{0}+v_{k}t,k})\,.$ (14)
By following the trajectories of the quasiparticles, we see that at time $t$ the quasiparticles that are shared between $A$ and its complement are the right-movers that originated at $x_{0}\in[-v_{k}t,\min(v_{k}t,\ell-v_{k}t)]$ and the left-movers that started at $x_{0}\in[\ell-\min(\left|v_{k}\right|t,\ell-\left|v_{k}\right|t),\ell+\left|v_{k}\right|t]$. Expressing everything in terms of the current position of the quasiparticle $x=x_{0}+v_{k}t$ for right-movers and $x=x_{0}-\left|v_{k}\right|t$ for left movers, we have
@@
To find the entanglement Hamiltonian of the mixed part of Eq. 2, we rewrite Eq. 14 as an exponential, using the property $\hat{n}_{x_{0},k}^{2}=\hat{n}_{x_{0},k}$. We obtain $\rho_{A}^{(t)}=e^{-K_{A,\text{ QP}}^{(t)}}/Z_{A}$, where $Z_{A}$ is a normalization and
@@
Analytic checks.
At sufficiently large times after the quench, the system relaxes locally to a stationary state which, generically for a free model is a generalized Gibbs ensemble (GGE) that incorporates all conserved charges. The Hilbert space of the pairs which are completely inside $A$ is empty and in the RDM (2) the pure part no longer contributes, while the EH (3) fully characterizes the RDM. In this limit, the time dependence of $\mathcal{K}_{L,R}$ drops out and asymptotic value of the EH (3) is
$=\int\frac{\mathrm{d}k}{2\pi}\,\eta(k)\,c_{k}^{\dagger}c_{k},$
The light-cone behavior of the EH and eventual relaxation is especially evident when the quenching Hamiltonian has a linear dispersion, $\varepsilon_{k}=vk$. In this case the quasiparticle velocity $v$ is independent of $k$ and the Heaviside functions can be taken outside of the mode integrals in (4), (5). Moreover, if $\eta(k)=\beta\varepsilon_{k}$, as happens after a quench in CFT [56], we find that
where $T$ and $\overline{T}$ are the right and left moving stress-energy operators. Eq. 20 agrees with previous results for the post-quench entanglement Hamiltonian in a CFT [32, 33]. Note that in this case the system relaxes locally to a Gibbs state in which only the Hamiltonian appears [57] and accordingly the EH has only short ranged hopping terms. This however, is not generically the case and $K_{A,\text{ QP}}^{(t)}$ can have a more intricate structure depending on both the initial state as encoded in $\eta(k)$ and $v_{k}$.
@@
$S^{(\alpha)}_{A}(t)=\int\frac{\mathrm{d}k}{2\pi}\,\min(2|v_{k}|t,\ell)\,h_{\alpha}(n(k))\,,$ (21)
@@
where $J_{\nu}(z)$ is Bessel’s function and we have introduced the asymptotic value of the correlation matrix $C_{i,j}^{(\infty)}=\delta_{i,j}/4+(\delta_{i,j-1}+\delta_{i,j+1})/2$. From the correlation matrix $C_{A}$ restricted to the subsystem $A$ it is possible to directly compute the entanglement Hamiltonian [60, 61, 62, 63]. Writing the lattice entanglement Hamiltonian as
@@
In order to compare the result of Eq. 24 with our prediction in Eq. 3, however, we need to take into account that at finite time Eq. 3 only describes the low lying part of the entanglement spectrum. Since according to Eq. 24 the higher part of the entanglement spectrum is given by the eigenvalues of $C_{A}$ that are close to either 00 or $1$, we introduce a cut-off on the spectrum, projecting out all the eigenvalues of Eq. 22 which are smaller than $10^{-4}$ or larger than $1-10^{-4}$. We achieve this by first diagonalizing $C_{A}$ then projecting on to the desired eigenspace after which we perform the inverse transformation and extract $h_{j,j+z}$ using (24).
@@
Conclusions and Outlook.
In this Letter, we derived a quasiparticle prediction, Eq. 3, for the EH after a global quantum quench in any system of non-interacting fermions. Remarkably, this simple form effectively captures the complex structure of the EH.
Our result paves the way for many intriguing generalizations. One key advantage of the QPP is its adaptability for calculating the entanglement entropy of disjoint intervals [36, 64], as well as more complex quantities like the negativity [41, 42]. Generalizing these results to an operator level, such as calculating the entanglement Hamiltonian of disjoint intervals or the negativity Hamiltonian [65, 66, 67], would be extremely interesting. Additionally, the QPP applies also for certain dissipative systems [68, 69, 70, 71], where also we could think of adapting our derivation. Furthermore, the QPP has been recently combined with dimensional reduction [72, 73] to describe the entanglement entropy [74] and asymmetry [75] in higher dimensional non-interacting models. The same could be done also for the EH. A significant challenge is generalizing our result to interacting integrable models, where the QPP breaks down in the calculation of Rényi entanglement entropy [76] and charged moments [45].
@@
Acknowledgments
We thank V. Alba, F. Ares, L. Capizzi, O. Castro-Alvaredo, B. Doyon, G. Di Giulio, J. Dubail, S. Murciano, and P. Zoller for fruitful discussions. All authors acknowledge support from ERC under Consolidator Grant number 771536 (NEMO). FR acknowledges support from the project “Artificially devised many-body quantum dynamics in low dimensions – ManyQLowD” funded by the MIUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 – Grant 2022R35ZBF.